40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 834     Entry time: Thu Aug 14 11:39:06 2008
Author: steve 
Type: Update 
Category: PEM 
Subject: particle counter is out of order 
The 40m ifo paricle counter sensor failed last night.
ELOG V3.1.3-